OpenStudy (anonymous):

Frances let a 17% tip on a meal, Including the tip, it's cost was 17.55. What was the cost of the meal before the tip.

OpenStudy (shubhamsrg):

let it be x now x+ (17% of x ) = x+ 0.17x = 17.55 just solve for x

OpenStudy (anonymous):

is the answer 20.55

OpenStudy (shubhamsrg):

nopes..please re-check..

OpenStudy (anonymous):

is the answer 14.55

OpenStudy (shubhamsrg):

it should be approximately 15.